# Molecular optomechanics with atomic antennas

### v1

@@ II.1 Modification of the properties of atomic antenna by a nanoparticle
    The plasmonic nanoparticle will modify the intensity of the incident laser at the site of the atomic antenna by a factor $K_{\text{inc}}^{2}$, enhancing its driving amplitude, and lowering the threshold for the excitation of the atomic antenna. To quantify this effect, we calculate the field distribution for illumination with a $p$-polarized plane wave incident at an angle of $60^{\circ}$, for a range of substrate-nanoparticle distances. We plot the field intensity normalized by the field intensity calculated in the absence of the nanoparticle (which already accounts for the imperfect transmission of the incident light into the substrate) in Fig. 2(a) with a solid blue line.
@@
    The plasmonic nanoparticle will also modify the intrinsic dissipation pathways for the atomic antenna, both by enhancing its coupling to the radiation modes, and by introducing nonradiative decay channels via absorption in the metal. The resulting decay rate of the atomic antenna will be increased, meaning that a larger laser intensity is required for its excitation. We model this effect by calculating the scattering from a dipolar electric source embedded in the diamond substrate, and oriented vertically, and quantify the enhancement of the decay rate by calculating the sum of the radiated and absorbed power ($P_{\text{rad}}+P_{\text{abs}}$), normalized by the emitted power in the absence of the nanoparticle $P_{0}$. Since this enhancement will only modify the rate of the radiative decay of the emitter $\gamma_{0}$, the overall decay rate can be estimated as
@@
    where we have again assumed that at the low temperature of 50 K, and in the absence of the nanoparticle, $\gamma=4\gamma_{0}$ [26]. We plot the resulting shortening of the antenna’s lifetime $(\gamma_{\text{tot}}/\gamma)^{-1}$ in Fig. 2(a) with a solid orange line  [^3] We note the limiting case of the decay rate enhancement from a buried electric dipole was considered in Ref. 50, where the enhancement in a similar system, with a plasmonic nanoparticle positioned on the dielectric substrate, exceeded $10^{3}$..
@@
    Finally, the presence of the nanoparticle will modify the field from the atomic antenna, experienced by the molecule. To quantify this effect, we calculate the electric field from the electric dipolar source positioned 2 nm below the surface, at the distance of 1 nm above the surface of the substrate, in the gap below the nanoparticle, and normalize it by the same quantity in the absence of the nanoparticle, to find the field enhancement factor $K_{\text{mol}}$ as a function of the substrate-nanoparticle spacing $d$. We plot the result in Fig. 2(a) with a dashed green line.
The above effects enter the conversion efficiency $\eta_{\sigma}$ by (i) boosting the effective coherent population $|\alpha_{\sigma}|^{2}$ by a factor given by the field intensity enhancement $K_{\text{inc}}^{2}$, (ii) increasing the decay rate $\gamma\rightarrow\gamma_{\text{tot}}$ and thus lowering $|\alpha_{\sigma}|^{2}$ by $\gamma/\gamma_{\text{tot}}$, and (iii) modifying the optomechanical coupling coefficient $g_{0,\sigma}$ by $K_{\text{mol}}$. In Fig. 2(b) we plot the resulting corrected conversion efficiency for the atomic antenna-enhanced Raman scattering,
@@
and the conventional plasmonic system with parameters given in Table 1. We find that while the nanoparticle can modify the efficiency of the atomic antenna-enhanced Raman scattering via the three effects we investigated, this overall change in efficiency shown in Fig, 2(b) is fairly small.
Figure 2: Effects of the nanoparticle (NP) on the dynamics of the atomic antenna system. In (a) we plot the enhancements of the intensity of: incident field $K_{\text{inc}}^{2}$ at the antenna position (solid blue line), field from the antenna at the molecules position (dashed green line), and the reduction of the lifetime $\gamma/\gamma_{\text{tot}}$ (solid orange line) of the antenna dipole. In (b) we plot the corrected conversion efficiencies for the atomic antenna (solid red line; Eq. (20)) and conventional plasmonic system (solid gray line; Eq. (18)).
@@
III Nonlinearities in atomic antenna-enhanced Raman scattering
The results discussed in the previous section were obtained within the linearized optomechanical picture, meaning that we assumed moderate laser intensities topping up to about $7\times 10^{-3}~{}\upmu\text{W}\cdot\upmu\text{m}^{-2}$ (see Table 1), at which the coherent amplitude of the atomic antenna peaks. Here we ask what happens beyond those intensities: does the simple linearisation scheme break down? Does the atomic antenna-enhanced Stokes emission exhibit a different dependence on laser intensity than a conventional plasmonic SERS setup?
Recall that in Subsection I.1, we linearized the optomechanical Hamiltonian by replacing the pseudospin operators with their expectation values. This approximation works well in the Heitler regime, where the coherent driving amplitude $\Omega$ of the atomic antenna is far smaller than its total decay rate $\gamma$. To better understand this effect, we used QuTip [45, 46] to model the dynamics of the system described by the full (un-linearized) optomechanical Hamiltonian ($\hat{H}_{0}+\hat{H}_{\text{coh}}+\hat{H}_{I}$) and the dissipative terms described in Subsection I.2. Given how weak are the effects discussed in Section II.1, for simplicity we use the uncorrected values of the parameters listed in Table 1.
In Fig. 3(a), we plot the calculated steady-state coherence ($|\langle\hat{\sigma}\rangle|^{2}$; solid blue line) and population ($\langle\hat{\sigma}^{\dagger}\hat{\sigma}\rangle$; dashed orange line) of the antenna, and find that for weak laser power $\langle\hat{\sigma}^{\dagger}\hat{\sigma}\rangle\approx|\langle\hat{\sigma}\rangle|^{2}$. As we increase the laser power towards the Mollow regime, the atomic antenna is driven into a maximally mixed state with $\left\langle\hat{\sigma}^{\dagger}\hat{\sigma}\right\rangle=0.5$. Furthermore, the incoherent emission spectrum of the antenna
@@ III Nonlinearities in atomic antenna-enhanced Raman scattering
develops the characteristic three-peak structure of a Mollow triplet (see Fig. 3(b)).
In that regime the system cannot absorb more energy, and we expect that the intensity of the Raman processes should saturate. Indeed, the Stokes (solid green line in Fig. 3(c)) and anti-Stokes (dashed red line) intensities, calculated from the incoherent spectrum of the plasmonic cavity (Eq. (I.4)), exhibits an initial linear dependence on $I$, followed by saturation.
Figure 3: Analysis of the (a) state and (b) emission from the atomic antenna, and (c) Stokes (solid green line) and anti-Stokes (dashed red line) intensities, as a function of intensity of the pump laser. Additionally, we plot the intensity of the Stokes emission in a conventional SERS setup (solid blue line). In (a) we observe a clear saturation of the low-intensity linear behaviour when the coherent pumping rate $\Omega$ approaches the total decay rate of the atomic defect $\gamma$, lowering the coherence $|\langle\hat{\sigma}\rangle|^{2}$ (blue) and saturating the population $\langle\hat{\sigma}^{\dagger}\hat{\sigma}\rangle$ (orange), as well as the Stokes emission intensity (c). Around this point the emission spectrum of the atom develops Mollow sidebands traced with the gray dashed lines. These results have been calculated using QuTiP, with parameters matching those listed in Table 1.
@@
In panel Fig. 3(c) we also plot, with solid blue line, the intensity of the Stokes emission from a conventional SERS setup, and find that below $I\sim 10^{-3}~{}\upmu\text{W}/\upmu\text{m}^{2}$ (the point of saturation of the atomic antenna-enhanced Stokes), this intensity is about three orders of magnitude weaker, consistent with the efficiency prediction shown in Fig. 2(b).
@@
We note that even for the largest pump intensity for which the linearization approximation would be justified (see the value in Table 1), the phonon populations do not diverge from the thermal equilibrium values ($n_{b}^{\mathrm{th}}\approx 5\times 10^{-2}$), and the optomechanical pumping and decay rates $\Gamma_{+}$ and $\Gamma_{-}$ are several orders of magnitude smaller than $\Gamma$ ($\Gamma_{+}/\Gamma\sim 10^{-9}\gg\Gamma_{-}/\Gamma$), suggesting that we would observe neither the vibrational pumping, nor instability.
Finally, we ask whether the nonclassical statistics of the electromagnetic field generated by the atomic antenna [47, 48] — in either Heitler or Mollow regimes — yields a similarly nonclassical statistics of the emitted Raman spectrum. We expect that should be the case, as the two effects that could normally suppress the sub-Poissonian statistics — incoherent pumping, and filtering by a spectrally narrower cavity — are absent. We have confirmed this by numerically calculating the intensity correlation $g^{(2)}(0)$ of the state of the plasmonic cavity mode $\hat{a}$, which mediates the scattering. We also performed calculations of the frequency-resolved intensity correlations between the Stokes lines, and Stokes-anti-Stokes pairs [15, 49]. These results are not shown here, as in each case we find perfect antibunching of the frequency-filtered emission.
IV Summary and outlook
@@
We identify two clear signatures of the atomic-antenna enhanced Raman scattering, that can be targeted for experimental verification: the first is the dependence of the intensity of Stokes emission on the laser intensity $I$, which illustrates a clear saturation, and a significant deviation from the conventionally observed in SERS spontaneous ($\propto I$), or vibrationally pumped ($\propto I^{2}$) emission. The second is the nonclassical statistics of the emission, due to the atomic antenna transducing the incident coherent light into a stream of antibunched photons. Additionally, the sub-GHz narrowband response of the atomic antenna can be utilized to identify and filter atomic antenna-enhanced Raman scattering. By comparing photoluminescence with on- and off- resonant laser excitation of the atomic antenna, background signals can be accurately identified, similar to observations of Förster resonant excitation transfer [26].
@@ Acknowledgements
Authors acknowledge fruitful discussions with Javier Aizpurua. MKS acknowledges funding from the Australian Research Council Discovery Early Career Researcher Award (DE220101272). AAH acknowledges funding from Q-NEXT, supported by the US Department of Energy, Office of Science, National Quantum Information Science Research Centers.
@@ Appendix A Derivation of the interaction Hamiltonian
The linearized optomechanical Hamiltonian reads
$\hat{H}_{I,\text{lin}}=-$ $\frac{1}{2}R_{ij}Q^{0}\left(\hat{b}+\hat{b}^{\dagger}\right)$
[…]
×$\times\left[(\mathbf{e}_{a})_{j}\hat{a}+(\mathbf{e}_{a})_{j}^{*}\hat{a}^{\dagger}\right].$ (22)

### v2

@@ II.2.1 Modification of the properties of atomic antenna by a nanoparticle
    The plasmonic nanoparticle will modify the intensity of the incident laser at the site of the atomic antenna by a factor $K_{\text{inc}}^{2}$, enhancing its driving amplitude, and lowering the threshold for the excitation of the atomic antenna. To quantify this effect, we calculate the field distribution for illumination with a $p$-polarized plane wave incident at an angle of $60^{\circ}$, for a range of substrate-nanoparticle distances. We plot the field intensity normalized by the field intensity calculated in the absence of the nanoparticle (which already accounts for the imperfect transmission of the incident light into the substrate) in Fig. 2(b) with a solid blue line.
@@
    The plasmonic nanoparticle will also modify the intrinsic dissipation pathways for the atomic antenna, both by enhancing its coupling to the radiation modes, and by introducing nonradiative decay associated with absorption in the metal. The resulting decay rate of the atomic antenna will be increased, meaning that a larger laser intensity is required for its excitation. We model this effect by calculating the scattering from a dipolar electric source embedded in the diamond substrate, and oriented vertically, and quantify the enhancement of the decay rate by calculating the sum of the radiated and absorbed power ($P_{\text{rad}}+P_{\text{abs}}$), normalized by the emitted power in the absence of the nanoparticle $P_{0}$. Since this enhancement will only modify the rate of the radiative decay of the emitter $\gamma_{0}$, (see the discussion in SI Section F.1) the overall decay rate can be estimated as
@@
    where we have again assumed that at the low temperature of 50 K, and in the absence of the nanoparticle, $\gamma=4\gamma_{0}$ [26]. We plot the resulting shortening of the antenna’s lifetime $(\gamma_{\text{tot}}/\gamma)^{-1}$ in Fig. 2(b) with a solid orange line. We note the limiting case of the decay rate enhancement from a buried electric dipole was considered in Ref. 47, where the enhancement in a similar system, with a plasmonic nanoparticle positioned on the dielectric substrate, exceeded $10^{3}$. Interestingly, due to the reflection from the interface between the diamond substrate and air, which modifies the radiative properties of the atomic antenna compared to its bulk response, this ratio does not converge to 1 in the limit of large spacings $d$.
@@
    Finally, the presence of the nanoparticle will modify the field from the atomic antenna, experienced by the molecule. To quantify this effect, we calculate the electric field from the electric dipolar source positioned 2 nm below the surface, at the distance of 1 nm above the surface of the substrate, in the gap below the nanoparticle, and normalize it by the same quantity in the absence of the nanoparticle, to find the field enhancement factor $K_{\text{mol}}$ as a function of the substrate-nanoparticle spacing $d$. We plot the result in Fig. 2(b) with a dashed green line.
The above effects enter the conversion efficiency $\eta_{\sigma}$ by (i) boosting the effective coherent population $|\alpha_{\sigma}|^{2}$ by a factor given by the field intensity enhancement $K_{\text{inc}}^{2}$, (ii) increasing the decay rate $\gamma\rightarrow\gamma_{\text{tot}}$ and thus lowering $|\alpha_{\sigma}|^{2}$ by $\gamma/\gamma_{\text{tot}}$, and (iii) modifying the optomechanical coupling coefficient $g_{0,\sigma}$ by $K_{\text{mol}}$. In Fig. 2(c) we plot the resulting corrected conversion efficiency for the atomic antenna-enhanced Raman scattering,
@@
and the conventional plasmonic system with parameters given in Table 1. We find that while the nanoparticle can modify the efficiency of the atomic antenna-enhanced Raman scattering via the three effects we investigated, this overall change in efficiency shown in Fig, 2(c) is within a factor of 4.
Figure 2: Effects of the nanoparticle (NP) on parameters of the atomic antenna setup depicted in the left panel of (a). (b) Enhancements of the intensity of: incident field $K_{\text{inc}}^{2}$ at the antenna position (solid blue line), field from the antenna at the molecules position (dashed green line), and the reduction of the lifetime $\gamma/\gamma_{\text{tot}}$ (solid orange line) of the antenna dipole. (c) Corrected conversion efficiencies for the atomic antenna (solid red line; Eq. (20)) and conventional plasmonic system (solid gray line; Eq. (18)).
@@
II.3 Nonlinearities in atomic antenna-enhanced Raman scattering
Previous sections explored the properties of Raman scattering in a linearized picture, which we derived in Subsection II.1.1 by replacing the pseudospin operator with its expectation value. This approximation works well in the Heitler regime, where the coherent driving amplitude $\Omega$ of the atomic antenna is far smaller than its total decay rate $\gamma$ [32, 48]. Here we ask what happens beyond this regime: does the linearization scheme break down? Does the atomic antenna-enhanced Stokes emission exhibit a different dependence on laser intensity than a conventional plasmonic SERS setup?
To answer these questions, we use QuTip [49, 50] to model the dynamics of the system described by the full (un-linearized) optomechanical Hamiltonian with additional coupling between the atomic and plasmonic antennas and conventional SERS coupling, and the dissipative terms described in Subsection II.1.2. Details of this model are given in Appendix F.
We start by analysing the response of the atomic antenna itself on the amplitude of the driving. First, in Fig. 3(a) we plot the incoherent emission spectrum of the antenna
@@ II.3 Nonlinearities in atomic antenna-enhanced Raman scattering
for increasing laser power. Colors of the spectra in Fig. 3(a) correspond to the laser intensities marked with vertical dashed lines in Fig. 3(b), where we plot the steady-state coherence of the antenna $|\langle\hat{\sigma}\rangle|^{2}$ (solid blue line) and the population of its excited state $\langle\hat{\sigma}^{\dagger}\hat{\sigma}\rangle$ (dashed orange line). For laser intensities beyond $I\sim 10^{-2}~{}\upmu\text{W}/\upmu\text{m}^{2}$ the antenna transitions from the Heitler to the Mollow regime, is driven towards the maximally mixed state with $\left\langle\hat{\sigma}^{\dagger}\hat{\sigma}\right\rangle=0.5$, and its emission spectrum develops the characteristic three-peak structure of a Mollow triplet.
In that regime the system cannot absorb more energy, and we expect that the intensity of the Raman processes should saturate or — due to the decrease of the coherence $\langle\hat{\sigma}\rangle$ and the effective optomechanical coupling — weaken. To test this, in Fig. 3(c) we plot the spectra of light emitted from the system for increasing laser intensities, which we again color-code to the vertical lines in Fig. 3(d). That panel traces the peak intensities of the Stokes (solid orange line) and anti-Stokes (solid blue line) emission, calculated from the incoherent spectrum of the plasmonic cavity (Eq. (II.1.4)).
This result clearly illustrates the breakdown of the linearization description, and suggests that the overall Raman response is governed by the population of the atomic antenna, whether it be coherent or incoherent. It is reminiscent of the features observed in other hybrid optomechanical systems, for example in resonant Raman scattering, where the vibronic coupling of the electronic transition (denoted by the pseudospin operator $\hat{\sigma}_{m}$, to differentiate it from the operator of the atomic antenna), and molecular vibrations of a molecule is governed by some version of the interaction Hamiltonian
$\hat{H}_{\text{res}}=g_{\text{res}}\hat{\sigma}_{m}^{\dagger}\hat{\sigma}_{m}(\hat{b}+\hat{b}^{\dagger}),$ (22)
accompanied by either a coherent driving of the electronic transition, either directly by the laser source, or mediated by a cavity mode. For example, in Ref. [51], Neuman and co-authors discuss the latter case, and find that Raman emission is proportional to the population of the excited electronic state $\left\langle\hat{\sigma}_{m}^{\dagger}\hat{\sigma}_{m}\right\rangle$, which has the same functional dependence as that discussed in Fig. 3. In Appendix C we analytically derive and analyze this correspondence in detail.
The interaction Hamiltonian given in Eq. (22) describes also a range of other hybrid mechanical systems, where a quantum dot [52, 53] or a superconducting circuit [54] is coupled to mechanical vibrations. In the former scenario, when the quantum dot resonantly couples to the optical field of a cavity, this interaction Hamiltonian can be transformed into the three-way coupling between the pseudo-spin, optical, and mechanical modes, similar to that given in Eq. (II.1.1).
@@
To complete the comparison with a conventional SERS setup, in panel Fig. 3(d) we also plot, with dashed orange and blue lines, the intensities of the Stokes and anti-Stokes emission from a plasmonic SERS system ($S_{a,\text{conv}}(\omega_{S}/\omega_{aS})$; calculated assuming vanishing $g_{0,\sigma}=0$). Below $I\sim 10^{-2}~{}\upmu\text{W}/\upmu\text{m}^{2}$ (the point of saturation of the atomic antenna-enhanced Stokes), its intensity is two to three orders of magnitude weaker than the antomic antenna-enhanced Raman scattering, consistent with the efficiency prediction shown in Fig. 2(c). Above $5~{}\upmu\text{W}/\upmu\text{m}^{2}$, the conventional SERS dominates the response.
Figure 3: Analysis of the (a) incoherent emission spectrum from the atomic antenna $S_{\sigma}$ (Eq. (21)) and (b) the coherence $|\langle\hat{\sigma}\rangle|^{2}$ (solid green) and population $\langle\hat{\sigma}^{\dagger}\hat{\sigma}\rangle$ (dashed orange) of the atomic antenna for increasing laser intensities. Colors of the spectra in (a) correspond to the vertical lines in (b) to denote intensities (explicitly listed in Appendix F). In (c) we plot the incoherent spectra of Raman emission $S_{a}$ (Eq. (10)) for increasing laser intensities; in (d) we trace the Stokes (solid orange) and anti-Stokes (solid blue) peak intensities, and compare them with the same calculated in conventional SERS setup (dashed lines). Details of this model and parameters are given in Appendix F.2 and in Table 1.
@@
Even for the largest pump intensity for which the linearization approximation would be justified (see the value in Table 1), the phonon populations do not diverge from the thermal equilibrium values ($n_{b}^{\mathrm{th}}\approx 5\times 10^{-2}$), and the optomechanical pumping and decay rates $\Gamma_{+}$ and $\Gamma_{-}$ are several orders of magnitude smaller than $\Gamma$ ($\Gamma_{+}/\Gamma\sim 10^{-9}\gg\Gamma_{-}/\Gamma$), suggesting that we would observe neither the vibrational pumping, nor instability.
Finally, we note that in the context of Surface Enhanced Resonant Raman Scattering (SERRS), Neuman and co-authors in Ref. [51] considered a system where the sidebands of Mollow triplet are frequency-matched to a Raman lines. Since the splitting between these sidebands is approximately given by coherent driving rate of the atomic antenna, such a setup would require an extremely strong coherent driving of the order of mechanical frequency $\sim$10-40 THz.
II.4 Feasibility
The experimental realization of the system presented in this work would be arguably challenging, since on top of the usual challenges of positioning the Raman-active molecule in the gap of a substrate-nanoparticle system, it adds the requirement of spatially aligning them with a shallow deposited GeV. This problem could be addressed by using sufficiently high density of defects and molecules, or by swapping the nanoparticle for a metallic tip in a Tip-Enhanced Raman Scattering experiment.
Using the estimates for $10^{-8}$ efficiency of thes Stoke emission, and considering laser intensity of $I\sim 10^{-2}~{}\upmu\text{W}/\upmu\text{m}^{2}$ (of flux of $3\times 10^{10}$ pump photons/s) which saturates the GeV, we can expect the Stokes flux to reach about 300 photons/s, consistent with the estimates for observation of Surface- and Tip-Enhanced Raman Scattering from single molecules [55, 56, 37, 5, 57].
III Summary and outlook
@@
We identify a clear signature of the atomic-antenna enhanced Raman scattering, that can be targeted for experimental verification — the strongly nonlinear dependence of the intensity of Stokes emission on the laser intensity $I$, which illustrates a clear saturation, and a significant deviation from the conventionally observed in SERS spontaneous ($\propto I$), or vibrationally pumped ($\propto I^{2}$) emission. Additionally, the sub-GHz narrowband response of the atomic antenna can be utilized to identify and filter atomic antenna-enhanced Raman scattering. By comparing photoluminescence with on- and off- resonant laser excitation of the atomic antenna, background signals can be accurately identified, similar to observations of Förster resonant excitation transfer [26].
@@ Acknowledgements
Authors acknowledge fruitful discussions with Javier Aizpurua and Ruben Esteban.
Funding sources
MKS acknowledges funding from the Australian Research Council Discovery Early Career Researcher Award (DE220101272). AAH acknowledges funding from Q-NEXT, supported by the US Department of Energy, Office of Science, National Quantum Information Science Research Centers.
@@ Appendix A Derivation of the interaction Hamiltonian
A.1 Complete Hamiltonian of the system
Figure 4: Schematics for the effects discussed in Section A.1.
In Fig. 4 we represent all the coherent processes that govern the dynamics of the system comprised of a single-mode plasmonic cavity, an atomic antenna modeled as a two-level system — both exposed to a coherent illumination by a laser — and a Raman-active molecule.
  1. 1.
    coherent driving of the atomic antenna
    $\hat{H}_{\text{coh}}=\hbar\Omega\left(e^{-i\omega_{l}t}\hat{\sigma}+e^{i\omega_{l}t}\hat{\sigma}^{\dagger}\right),$ (23)
    and of the optical cavity mode
    $\hat{H}_{\text{coh},a}=\hbar\Omega_{a}\left(e^{-i\omega_{l}t}\hat{a}+e^{i\omega_{l}t}\hat{a}^{\dagger}\right);$ (24)
    The two amplitudes $\Omega$ and $\Omega_{a}$ can be derived from the laser intensity $I$ and the system parameters, as discussed in Appendix E and Section II.1.1; throughout the manuscript, we assume laser to be tuned to the dominant transition in GeV,
  2. 2.
    the direct coupling between the optical cavity and the atomic antenna $\hat{H}_{a-\text{GeV}}$
    $\hat{H}_{a-\text{GeV}}=-\hat{\mathbf{E}}_{a}(\mathbf{r}_{\sigma})\cdot\mathbf{d}_{\sigma}\approx\hbar g_{JC}(\hat{a}\hat{\sigma}^{\dagger}+\hat{a}^{\dagger}\hat{\sigma}),$ (25)
    where we performed the rotating wave approximation.
    In Section II.1, we assume that this resonant coupling will result solely in the enhancement of the emission rate from the molecule via the Purcell effect. The magnitude of this effect for a system with parameters described Table 1 can be calculated numerically. From the numerical simulations, the maximum Purcell effect from the plasmonic antenna (see the discussion in Section II.2.1 and Section F.1) can be estimated as $F_{P}\approx 30$. If we assume that this effect is due to the coupling with a detuned single optical mode, $F_{P}$ should be given as
    $F_{P}=\frac{4g_{JC}^{2}\kappa}{[\kappa^{2}+(\omega_{\sigma}-\omega_{a})^{2}]\gamma},$ (26)
    giving $g_{JC}/2\pi\approx 0.38$ THz. For larger spacing, this would drop to $0.1$ THz.
    These estimates place the effect deep in the weak coupling regime. If we nevertheless wanted to diagonalise the system made up of the optical cavity and atomic antenna, the mixing angle would be $g_{JC}/(\omega_{\sigma}-\omega_{a})$ would be of the order of 0.01 (assuming detunings as discussed in the main text), justifying the perturbative treatment of the optomechanical coupling.
  3. 3.
    interaction between the Raman dipole and the electric fields of the optical cavity, and the atomic antenna $\hat{H}_{I}$
    $\hat{H}_{I}=-\frac{1}{2}\hat{\mathbf{p}}_{R}\cdot\left[\hat{\mathbf{E}}_{a}(\mathbf{r}_{m})+\hat{\mathbf{E}}_{\sigma}(\mathbf{r}_{m})\right].$ (27)
    The Raman dipole $\hat{\mathbf{p}}_{R}$ is induced by both the electric fields of the optical cavity and the atomic antenna oscillating at the pump frequency $\omega_{l}$, as $\hat{\mathbf{p}}_{R}=R:\left[\hat{\mathbf{E}}_{a}(\mathbf{r}_{m})+\hat{\mathbf{E}}_{\sigma}(\mathbf{r}_{m})\right]$, and the above interaction Hamiltonian can be expanded as
    $\hat{H}_{I}=-\frac{1}{2}\hat{x}\bigg{[}$ $(R:\hat{\mathbf{E}}_{a})\cdot\hat{\mathbf{E}}_{a}$
    $+(R:\hat{\mathbf{E}}_{a})\cdot\hat{\mathbf{E}}_{\sigma}$
    $+(R:\hat{\mathbf{E}}_{\sigma})\cdot\hat{\mathbf{E}}_{a}$
    $\left.+(R:\hat{\mathbf{E}}_{\sigma})\cdot\hat{\mathbf{E}}_{\sigma}\right].$ (28)
    The first term describes the conventional optomechanical coupling, and processes where the cavity photons drive the Raman processes in the molecule, and the scattered Stokes and anti-Stokes photons couple into the cavity. This mechanism, together with the coherent pumping term $H_{\text{coh},a}$, will be referred to as the conventional plasmonic SERS. Parameters of this setup are discussed in Section E.
    The second line describes a similar processes, but with the Raman emission coupled to the atomic antenna, instead of the optical cavity mode. Since the pump laser is tuned to the dominant transition in GeV at 602 nm, and linewidth of this transition is several orders of magnitude smaller than the molecular vibration, this process cannot be frequency-matched.  [^1] Once could imagine a different scenario, where the pump was tuned to match a particular Stokes or anti-Stokes line to the transition in GeV. However, since the linewidth of the GeV transition is far smaller than the typical decay rate of a molecular vibration, the former could not be treated as a bath, complicating the analytical treatment.
    The third line describes the effect denoted schematically in Fig. 1, where the incident illumination is resonantly scattered on the atomic antenna, drives the Raman scattering on the molecule, and the Raman photons couple to the spectrally broad optical cavity mode. We focus on this mechanism throughout this work.
    Finally, in the last line we consider an extremely weak effect, where the Raman dipole is induced by light resonantly scattered on the atomic antenna, and the Raman emission coupling back to the antenna. Since the linewidth of the transition in GeV is about 5 orders of magnitude smaller than the vibrational frequency, necessarily either one of these processes will be off-resonant and thus its contribution will be negligible.
A.2 Derivation of the linearized interaction Hamiltonian
We now focus on the atomic antenna-enhanced Raman scattering. The linearized optomechanical Hamiltonian reads
$\hat{H}_{I,\text{lin}}=-$ $\frac{1}{2}R_{ij}Q^{0}\left(\hat{b}+\hat{b}^{\dagger}\right)\left[\omega_{\sigma}^{2}\mu_{0}(\mathbf{G}(\mathbf{r}_{\sigma},\mathbf{r}_{m},\omega_{\sigma})\cdot\mathbf{d}_{\sigma})_{i}\alpha_{\sigma}+\text{h.c.}\right]\left[(\mathbf{e}_{a})_{j}\hat{a}+(\mathbf{e}_{a})_{j}^{*}\hat{a}^{\dagger}\right].$ (29)